OpenStudy (anonymous):

why does 6 to the zero power equal 1 explain

OpenStudy (john_es):

\[6^0=6^{(+1-1)}=6^1\cdot6^{-1}=\frac{6}{6}=1\]

OpenStudy (anonymous):

ok i dont get that but thanks